OpenStudy (anonymous):

How to solve: Sin(70)Cos(10) - Sin(10)Cos(70) Can you simplify to Sin(60)Cos(-60) or can you not do that?

jimthompson5910 (jim_thompson5910):

Sin(70)Cos(10) - Sin(10)Cos(70) is the same as Sin(70)Cos(10) - Cos(70)Sin(10) ------------------------------------------------------- use sin(A-B) = sin(A)*cos(B) - cos(A)*sin(B)

OpenStudy (anonymous):

So I get Sin(70-10) = Sin70Cos10 - Cos70Sin10 Sin (70-10) = Sin(60) = Sq. 3/2 Did I do that right?

jimthompson5910 (jim_thompson5910):

you did that perfectly, the answer is \[\Large \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Thank you once again!

jimthompson5910 (jim_thompson5910):

sure thing